IhelpUuHELPme:

@jhonyy9 @MiraAngel For this question, an answer that is not a whole number should be entered as a fraction in simplest form using / as the fraction bar. Wendy is creating a large soup based on a recipe. Her recipe calls for 3/4 of a pint of milk, but she only has 5/8 of a pint of milk, so she can only make a portion of her soup. Wendy can only make _______ of her recipe with the amount of milk that she has

IhelpUuHELPme:

there is no diagram

darkknight:

basically she needs 3/4, but she only has 5/8. What fraction of 5/8 is 3/4?

darkknight:

So you want to do 5/8 divided by 3/4 and remember that dividing is the same as multiplying by the reciprocal, so do 5/8 times 4/3, what you get is how much of the recipe wendy can make with the amount of milk she has

snowflake0531:

\[\frac{ 5 }{ 8 } \div \frac{ 3 }{ 4 } = \frac{ 5 }{ 8 } \times \frac{ 4 }{ 3 } = \frac{ 5 }{ 2 } \times \frac{ 1 }{ 3 } = \frac{ 5 }{ 6 }\]

snowflake0531:

😔🤧😔🤧😔🤧

IhelpUuHELPme:

quit it .__.